OpenStudy (anonymous):

solve (1 + x^2) y'' -2 xy' +3y = e^-x using power of series.

OpenStudy (amistre64):

what have you worked out so far?

OpenStudy (anonymous):

I am confused about the taylor series for e^(-x)

OpenStudy (amistre64):

have you already plugged in your y, y' and y'' series the rest is just comparison right

OpenStudy (anonymous):

yes i have all that. how can i compare to its taylor series(which i dont know) but i do know for e^x

OpenStudy (amistre64):

then let me catch up \[(1 + x^2) y'' -2 xy' +3y = e^{-x}\] \[(1 + x^2) \sum_2a_n~n(n-1)x^{n-2} -2 x\sum_1a_n~nx^{n-1} +3\sum_0a_nx^n = e^{-x}\] \[\sum_2a_n~n(n-1)x^{n-2} + \sum_2a_n~n(n-1)x^{n} -2 \sum_1a_n~nx^{n} +3\sum_0a_nx^n = e^{-x}\] \[\sum_0a_{n+2}~(n+2)(n+1)x^{n} + \sum_2a_n~n(n-1)x^{n} -2 \sum_1a_n~nx^{n} +3\sum_0a_nx^n = e^{-x}\] \[(3a_0+2a_{2})+(5a_1+6a_{3})x+\\ ~~~~ \sum_2 [a_{n+2}~(n+2)(n+1) +[a_n~n(n-1)-2a_n~n+3a_n]~x^n = e^{-x}\] so far so good? or have a messed it up along the way

OpenStudy (amistre64):

\[e^u = \sum_0 \frac{1}{n!}u^n\] \[e^{-x} = \sum_0 \frac{1}{n!}(-x)^n\] \[e^{-x} = \sum_0 \frac{(-1)^nx^n}{n!}\]

OpenStudy (amistre64):

\[e^{-x}=\frac{1}{0!}-\frac{1}{1!}x+\sum_2\frac{(-1)^nx^n}{n!}\]

OpenStudy (anonymous):

On your last equation above, isnt that supposed to be \[(6a_3+a_1)\]

OpenStudy (amistre64):

maybe, its a bit difficult to keep track of when coding :)

OpenStudy (anonymous):

it's okay. so now then how do we compare?

OpenStudy (amistre64):

\[ a_{0+2}~(0+2)(0+1)x^{0} \\ +a_{1+2}~(1+2)(1+1)x^{1} \\ - 2a_1~1x^{1} \\ +3a_0x^0 \\ + 3a_1x^1 \\ \] \[ 2a_{2} +6a_{3}~x - 2a_1~x +3a_0 + 3a_1x \]

OpenStudy (amistre64):

well, as long as our first 2 terms are fit to the first 2 terms of e^-x the rest are recurrsive

OpenStudy (amistre64):

1 = (3a0 +2a2) -1x = (6a3+a1)x \[a_{n+2}~(n+2)(n+1)+a_n~n(n-1)-2a_n~n+3a_n = \frac{(-1)^n}{n!}~:~n\ge2\] solve for a_n or a_(n+2)

OpenStudy (amistre64):

looks like we will be able to define everything we need to in terms of a_0 and a_1

OpenStudy (anonymous):

okay this does make sense to me now. i was confused about the taylor series part Thanks alot for your help

OpenStudy (amistre64):

good luck :)